# TOGLL: Correct and Strong Test Oracle Generation with LLMs

### v1

@@ III-E Threats to Validity
In our investigation, during fine-tuning we utilized the ‘exact match’ metric to compute oracle generation accuracy. This is a widely used metric by the research community. For unseen inference, we relied on test validation. This methodology, we assert, offers a reliable accuracy metric. Additionally, for measuring test oracle strength, we used mutant kill score, a well established metric used in many research studies  [51, 16, 15].
@@ IV-B Token Length Limits
We have seen that LLMs can not process input samples that exceeds the maximum token length capacity of the LLMs. In our study, a total of 3% of the samples exceeded the threshold of 600 tokens when using prompt 6 that includes MUT, docstring and prefix. Rather than increasing the token threshold, our study suggests that in such cases we could shift to a more concise prompt, such as prompt 3 which replaces the MUT with the method signature without significantly compromising performance.
@@ V Related Work
In the past year, techniques for test generation using LLMs have emerged. For example, TESTPILOT  [42], an LLM-based test generation tool for JavaScript that automatically generates unit tests for JavaScript. TESTPILOT achieved 70% statement coverage and 52.8% branch coverage, which outperformed the state-of-the feedback-directed JavaScript test generation technique, Nessie [2]. A recent study  [44] that utilized LLMa for generating test for the SF110 large-scale real world Java projects showed that LLMa could only achieve 2% coverage, whereas EvoSuite can achieve more than 90% coverage. This suggests that LLMs are not very effective in generating test for real world Java program, which is why in our study instead of focusing on both test input and test oracle generation, we focus on a single task test oracle generation with LLM.
@@ VI Conclusion
We conduct the first large-scale investigation of the ability of large language models to automatically generating correct and strong test oracles. To this end, we fine-tune seven code LLMs that are pre-trained to perform code related tasks. We fine-tune CodeGPT-110M, CodeParrot-110M, CodeGen-350M, PolyCoder-.4B, Phi-1(1.3B), CodeGen-2B and PolyCoder-2.7B on six different prompts with varying level of information. Our findings reveal that effective prompt design can improve the accuracy of the models significantly and that with effective prompts smaller size model can also perform equally or even better than larger models. As fine-tuning and inference with large models is environmentally and financially costly [4], small models are most preferred.
With the best performing model and prompt we develop TOGLL, our LLM-based automated test oracle generation method. We extensively evaluate TOGLL through large-scale study to investigate its generalizability and capability to generate correct, diverse and strong test oracles. Our experimental study answers critical research questions, demonstrating that TOGLL can generate 3.8x more correct assertion and 4.9x more exception oracles than the previous SOTA neural method, TOGA. Not only does TOGLL significantly reduce the generation of false positive oracles, but it generates oracles 97% of the time in comparison to TOGA which only generates oracles 38% of the time. Generating correct oracle does not necessarily mean that the generated assertions are diverse and strong enough to detect faults. Our results indicate that TOGLL is capable of generating diverse test oracle where only 9.5% are an exact match with the unseen dataset. Furthermore, TOGLL’s diverse assertions can detect 1,023 unique bugs that EvoSuite can not detect with it own generated assertions, which is 10x improvement over prior ML oracle generation techniques. TOGLL establishes a new state-of-the-art for ML-based oracle generation.
We describe remaining challenges in using LLMs for test oracle generation and discuss several actionable future directions to mitigate these limitation. We believe our large-scale study and in-depth investigation lay down the groundwork for further advancements in automated test oracle generation.

### v2

@@ III-F Threats to Validity
In our investigation, during fine-tuning we utilized the ‘exact match’ metric to compute oracle generation accuracy. This is a widely used metric by the research community. For unseen inference, we relied on test validation. This methodology, we assert, offers a reliable accuracy metric. Additionally, to complement the results from mutation testing, we performed a real bug detection study. Both metrics are well-established and used in many research studies [7, 5, 4].
@@ IV-B Token Length Limits
We have seen that LLMs can not process input samples that exceeds the maximum token length capacity of the LLMs. In our study, a total of 3% of the samples exceeded the threshold of 600 tokens when using P6 that includes MUT, docstring and prefix. Rather than increasing the token threshold, our study suggests that in such cases we could shift to a more concise prompt, such as P3 which replaces the MUT with the method signature without significantly compromising performance.
@@ V Related Work
In the past year, techniques for test generation using LLMs have emerged. For example, TESTPILOT  [21], an LLM-based test generation tool for JavaScript that automatically generates unit tests for JavaScript. TESTPILOT achieved 70% statement coverage and 52.8% branch coverage, which outperformed the state-of-the feedback-directed JavaScript test generation technique, Nessie [53]. A recent study  [54] that utilized LLMs for generating test for the SF110 large-scale real world Java projects showed that LLMa could only achieve 2% coverage, whereas EvoSuite can achieve more than 90% coverage. This suggests that LLMs are not very effective in generating test for real world Java program, which is why in our study instead of focusing on both test input and test oracle generation, we focus on a single task test oracle generation with LLM.
@@ VI Conclusion
We conduct the first large-scale investigation of the ability of large language models to automatically generating correct and strong test oracles. To this end, we fine-tune seven code LLMs with six different prompts with varying level of information. Our findings reveal that effective prompt design can improve the accuracy of the models significantly and that with effective prompts smaller size model can also perform equally or even better than larger models.
With the best performing model and prompt we develop TOGLL, our LLM-based automated test oracle generation method. We extensively evaluate TOGLL through large-scale study to investigate its generalizability and capability to generate correct, diverse and strong test oracles. Our experimental study answers critical research questions, demonstrating that TOGLL can generate 3.8x more correct assertion and 4.9x more exception oracles than the previous SOTA neural method, TOGA. Generating correct oracle does not necessarily mean they are diverse and strong enough to detect bugs [55]. We find that TOGLL is capable of generating diverse test oracle where only 9.5% are an exact match with the ground truth. Furthermore, TOGLL’s diverse assertions can detect a large number of unique bugs that EvoSuite can not detect. For both mutant and real bug detection, TOGLL significantly outperformed TOGA, establishing itself as the new SOTA.
We describe remaining challenges and discuss several actionable future directions. We believe our large-scale study and in-depth investigation lay down the groundwork for further advancements in automated test oracle generation.
VII Acknowledgment
This material is based in part upon work supported by National Science Foundation awards 2129824 and 221707. The authors acknowledge Research Computing at The University of Virginia for providing computational resources and technical support that have contributed to the results reported within this publication.